# Unified Modeling and Structural Optimization of Multi-magnet Embedded Soft Continuum Robots for Enhanced Kinematic Performances

### v1

@@ III-B Optimization Methodology
Figure 7: a.) Optimization results under $2^{2}$ magnetic orientation combinations. The background color maps represent normalized performance values obtained via exhaustive search. Black lines indicate trajectories of the gradient-based optimization. b.) Optimization results across $2^{3}$ magnetic orientations, comparing the gradient-based method with a heuristic method. Bar plots show the configuration at optimized manipulability with convergence curves overlaid.
III-B1 Analytical Example
Consider a MeSCR with two axially embedded magnets actuated by a spatially uniform field $\mathbf{b}_{u}\in\Pi$ with $\mathbf{v}^{\top}\bar{\mathbf{t}}=0$, where the global manipulability is to be maximized [23]. The two magnets have identical dipole strength. Let one magnet be fixed at the distal end ($L_{k_{1}}=L$), while the position of the other magnet $L_{k_{0}}$ serves as the design variable, subject to the boundary constraint $0<L_{k_{0}}<L$. From Corollary 1, the joint variable simplifies to $\boldsymbol{\theta}_{i}=\theta_{i}\mathbf{n}$. In this configuration, the MeSCR becomes kinematically equivalent to a planar $N$-joint revolute open-chain mechanism. Let $\theta_{i}^{j}\triangleq\sum_{l=i}^{j}\theta_{l}$ denote the cumulative joint angle. In the regime of weak magnetic actuation, the global manipulability index has the linearized expression on $\theta_{k_{0}}^{k_{1}}$: $z\approx L_{k_{0}}(L-L_{k_{0}})\left|\theta_{k_{0}}^{k_{1}}\right|$, where the analytical form is derived from the planar 2-joint revolute open-chain mechanism. Similarly, applying the approximate $\mathbf{G}_{m}$ from (23), the immersion Jacobian also admits the linearized form:
$\mathcal{J}_{\boldsymbol{\theta}}\approx$
$\begin{cases}\!M^{3}\!\sqrt{\Lambda_{k_{0}}^{m}(\Lambda_{k_{1}}^{m}-\Lambda_{k_{0}}^{m})(3\Lambda_{k_{0}}^{m}+\Lambda_{k_{1}}^{m})}\left|\theta_{k_{0}}^{k_{1}}\!\right|,\text{if }\bar{\mathbf{m}}_{k_{0}}\!=\!\bar{\mathbf{m}}_{k_{1}},\\ \!M^{3}\!\sqrt{\Lambda_{k_{0}}^{m}}(\Lambda_{k_{1}}^{m}-\Lambda_{k_{0}}^{m})\left|\theta_{k_{0}}^{k_{1}}\!\right|,\text{if }\bar{\mathbf{m}}_{k_{0}}=-\bar{\mathbf{m}}_{k_{1}},\end{cases}$
The cumulative rotation angle can also be approximated using (17) as $\left|\theta_{k_{0}}^{k_{1}}\right|\approx\frac{L-L_{k_{0}}}{EI}\left|\bar{\mathbf{m}}_{k}^{\top}[\mathbf{b}_{u}]_{\times}\mathbf{n}\right|$. Substituting this into the above expressions for $z$ and $\mathcal{J}_{\boldsymbol{\theta}}$, we can factor the dependence on $L_{k_{0}}$ out of the integral in the objective. The structural optimization problem thus reduces to:
$\max_{L_{k_{0}}}\begin{cases}\!L_{k_{0}}^{2}(L-L_{k_{0}})^{4}\sqrt{4L_{k_{0}}^{2}\!+\!(L-L_{k_{0}})^{2}},\mbox{if}\ \bar{\mathbf{m}}_{k_{0}}\!=\!\bar{\mathbf{m}}_{k_{1}},\\ \!L_{k_{0}}^{2}(L-L_{k_{0}})^{5},\mbox{if}\ \bar{\mathbf{m}}_{k_{0}}\!=\!-\bar{\mathbf{m}}_{k_{1}}.\end{cases}$
$L_{k_{0}}^{\ast}\approx\begin{cases}0.368L,\ \text{if }\bar{\mathbf{m}}_{k_{0}}=\bar{\mathbf{m}}_{k_{1}},\\ \frac{2}{7}L,\ \text{if }\bar{\mathbf{m}}_{k_{0}}=-\bar{\mathbf{m}}_{k_{1}}.\end{cases}$
We also evaluated the kinematic dexterity of the MeSCR through minimizing the distortion as defined in [24]. It is easy to verify that the distortion density satisfies the assumption of Theorem 5. Thus, we deduce that the optimal configuration of embedded magnets has a boundary distribution. We show numerical simulation results of normalized manipulability and dexterity in Fig. 6a When the magnetic moments are aligned (left graph), the configuration that maximizes manipulability exhibits good agreement with the theoretical result. In contrast, the opposing case (right graph) shows a deviation, which may be attributed to stronger nonlinearity in the index. For both cases, the configurations that minimize distortion consistently place the embedded magnet at the distal end, as expected. Fig. 6b illustrates the reachable distal trajectories under varying magnet placements. Although the optimal placements yield high manipulability along the path, their spatial extent remains limited. This apparent disparity underscores the geometric distinction between local controllability and global reachability, while the latter pertains to the image of the equilibrium map over the actuation submanifold. Therefore, any attempt to maximize overall performance must reconcile the trade-off between local dexterity and global coverage, as the two are not generally aligned.
III-B2 Numerical Examples
For MeSCRs with three or more embedded magnets, deriving a closed-form solution to the optimization problem becomes increasingly intractable. In such cases, we resort to a gradient-based optimization scheme, leveraging the analytical gradient expression provided in (22). Fig. 7a illustrates the optimization trajectories for representative cases with four possible direction combinations. The background colormaps depict the globally normalized performance values obtained via exhaustive enumeration. Across all cases, the gradient-based method consistently converges to configurations that match the global optima, demonstrating its reliability and consistency with full search results.
Moreover, the optimized magnet placements for maximizing dexterity closely follow the geometric insights established in Theorem 5. In all observed cases, the optimizer implicitly minimizes the net torque induced by the embedded magnets. Specifically, when adjacent magnets are aligned, the solution tends to separate them spatially to avoid torque amplification, whereas when they are oppositely oriented, the optimizer brings them closer to enable partial cancellation. These outcomes suggest that the dexterity metric inherently penalizes unbalanced torque distributions that restrict the range of achievable actuation directions.
As shown in Fig. 7b, the heuristic and gradient-based optimization methods yield consistent results both in terms of optimal configuration and convergence behaviors. Combining with the performance maps in Fig. 7a, it demonstrates that manipulability is highly sensitive to the spatial layout and directions of the embedded magnets. In some cases, optimal configurations are observed when magnets are distributed farther apart, highlighting the benefit of spatially diverse actuation sources. In comparison to dexterity, manipulability optimization focuses on maximizing the volume of attainable motion directions. As a result, the optimizer may favor configurations that introduce local torque in exchange for broader kinematic responsiveness.
IV Conclusion and Discussion
This paper presents a unified framework for modeling and optimizing the kinematic performance of multi-magnet embedded soft continuum robots. The proposed EPRB model enables a differentiable representation of equilibrium configurations under magnetic actuation, and supports explicit expressions for both the manipulator Jacobian and Hessian in multi-magnet settings. The analysis of equilibrium well-posedness yields sufficient conditions for solution uniqueness and continuity under bounded actuation, offering a means to ensure continuous configuration tracking in quasi-static operations. The upper bound on controllable degrees of freedom, shown to be twice the number of embedded magnets, offers a simple yet practical rule for relating design parameters to task-space controllability.
A key contribution of this paper lies in the development of a differential geometric optimization framework, where classical kinematic measures are reparametrized over the actuation space via equilibrium immersions. This approach accommodates a wide range of performance criteria, including but not limited to manipulability and dexterity, and admits both analytic and gradient-based solution methods. Although demonstrated in the context of MeSCRs, the formulation applies more broadly to continuum structures actuated by distributed torques, and may offer guidance in the design of other soft or compliant mechanisms where actuation and configuration are implicitly coupled.
Appendix A Proofs of Propositions
A-A Proof of Proposition 1
Noticing that
$\|\mathbf{M}_{i}(\boldsymbol{\theta})\mathbf{b}\|$ $=\left\|\sum_{k\in\mathbb{K},k\geq i}\left(\left[\mathbf{m}_{k}\right]_{\times}^{\top}\mathbf{R}_{0}^{i}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}\right)^{\top}\mathbf{b}(\mathbf{p}_{k})\right\|$
$\leq\left\|\mathbf{R}_{0}^{i}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}\right\|\sum_{k\in\mathbb{K},k\geq i}\left\|\left[\mathbf{m}_{k}\right]_{\times}\mathbf{b}(\mathbf{p}_{k})\right\|$
$=\sum_{k\in\mathbb{K},k\geq i}M_{k}B_{k}\triangleq\mathcal{M}_{i},$
it is convenient to obtain:
$\|\mathbf{M}(\boldsymbol{\theta})\mathbf{b}\|$ $=\sqrt{\sum_{i=0}^{N-1}\|\mathbf{M}_{i}(\boldsymbol{\theta})\mathbf{b}\|^{2}}\leq\sqrt{\sum_{i=0}^{N-1}\mathcal{M}_{i}^{2}}\triangleq\mathcal{M},$
and
$\|\mathbf{M}(\boldsymbol{\theta})\|$ $=\sqrt{\sum_{i=0}^{N-1}\|\mathbf{M}_{i}(\boldsymbol{\theta})\|}\leq\sqrt{N_{m}k_{N_{m}-1}}\max_{k}M_{k}\triangleq\mathcal{M}_{0}.$
A-B Proof of Proposition 2
Given that the Hessian matrix is symmetric, it can be expressed equivalently as
$\frac{\partial^{2}\mathcal{E}_{m}}{\partial\boldsymbol{\theta}\partial\boldsymbol{\theta}^{\top}}=\sum_{k\in\mathbb{K}}\operatorname{sym}\left(\mathbf{D}_{k}^{\top}\mathbf{C}_{k}\mathbf{D}_{k}+\mathbf{Q}^{k}\right)\triangleq\mathbf{S}_{m}.$
Whence,
$\mathbf{D}_{k}=\operatorname{blkdiag}\{\mathbf{R}_{0}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{0}},\dots,\mathbf{R}_{0}^{k}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{k}},\underbrace{\mathbf{0}_{3\times 3},\dots,\mathbf{0}_{3\times 3}}_{N-k-1}\},$ (24)
$\mathbf{Q}^{k}=\operatorname{blkdiag}\{\mathbf{Q}_{0},\dots,\mathbf{Q}_{k},\underbrace{\mathbf{0}_{3\times 3},\dots,\mathbf{0}_{3\times 3}}_{N-k-1}\},$
and
$\mathbf{C}_{k}=\operatorname{blkdiag}\{2\mathbf{U}_{k\times k}-\mathbf{I}_{k\times k},\underbrace{0\dots 0}_{N-k-1}\}\otimes\left([\mathbf{b}_{k}]_{\times}^{\top}[\mathbf{m}_{k}]_{\times}\right)$
with $\mathbf{U}_{k\times k}$ denotes the upper triangle matrix with all elements equal to $1$ and $\otimes$ denotes the Kronecker product. Then, using the triangle inequality and sub-multiplicative property, the following inequality holds for the spectrum norm of the Hessian matrix:
$\left\|\mathbf{S}_{m}\right\|$ $=\left\|\sum_{k\in\mathbb{K}}\operatorname{sym}\left\{\mathbf{D}_{k}^{\top}\mathbf{C}_{k}\mathbf{D}_{k}+\mathbf{Q}^{k}\right\}\right\|$ (25)
$\leq\sum_{k\in\mathbb{K}}\left\|\mathbf{D}_{k}^{\top}\right\|\left\|\mathbf{C}_{k}\right\|\left\|\mathbf{D}_{k}\right\|+\left\|\mathbf{Q}^{k}\right\|$
$=\sum_{k\in\mathbb{K}}\left\|\mathbf{C}_{k}\right\|+\left\|\mathbf{Q}^{k}\right\|$
where we used the fact that $\left\|\mathbf{D}_{k}^{\top}\right\|=\left\|\mathbf{D}_{k}\right\|=1$. The spectrum norm of $\mathbf{C}_{k}$ is straightforward to be given by
$\left\|\mathbf{C}_{k}\right\|$ $=\left\|2\mathbf{U}_{k\times k}-\mathbf{I}_{k\times k}\right\|\left\|[\mathbf{b}_{k}]_{\times}^{\top}[\mathbf{m}_{k}]_{\times}\right\|$ (26)
$\leq\frac{4k}{\pi}\|\mathbf{b}_{k}\|\|\bar{\mathbf{m}}_{k}\|=\frac{4k}{\pi}M_{k}B_{k}$
For small rotation angles, we note that
$\lim_{\boldsymbol{\theta}\to\mathbf{0}}\left\|\mathbf{Q}^{k}\right\|$ $=\lim_{\boldsymbol{\theta}\to\mathbf{0}}\max_{i=0,\dots,k}\left\{\left\|\mathbf{Q}_{i}\right\|\right\}=0.$ (27)
Combing (25), (26), and (27), the Jacobian of the vector-valued function $\mathbf{M}(\boldsymbol{\theta})\mathbf{b}$ satisfies:
$\left\|\frac{\partial}{\partial\boldsymbol{\theta}}\mathbf{M}(\boldsymbol{\theta})\mathbf{b}\right\|=\left\|-\mathbf{S}_{m}\right\|\leq\sum_{k\in\mathbb{K}}\frac{4k}{\pi}M_{k}B_{k}\triangleq\mathcal{L}.$
Therefore, the function is Lipschitz continuous on $\mathbb{R}^{3N}$ with the Lipschitz constant $\mathcal{L}$.

### v2

@@ IV-C2 Optimal Placement
We also evaluated the kinematic dexterity of the MeSCR by minimizing the distortion index in [22]. Since the distortion density satisfies the conditions of Theorem 4, the optimal solutions are expected to concentrate near boundary configurations. The numerical results in Fig. 2 confirm this behavior. Across both aligned and opposing magnetization modes, the distortion minimizing designs consistently collapse toward the distal end, whereas the manipulability maximizing designs follow the placement trends predicted by the analytical model. These outcomes highlight a clear structural organization of the performance landscape. Manipulability is governed by the separation and orientation of magnetic sources, while distortion tends to favor extreme placements that suppress curvature variation. The reachable trajectories further show that improved local dexterity does not necessarily translate into expanded global coverage. This observation reinforces the intrinsic mismatch between local controllability and global reachability and indicates that structural optimization must balance these competing geometric effects rather than optimizing either in isolation.
IV-D Numerical Results for Multi-magnet Configurations
For MeSCRs with three embedded magnets, closed-form optimization becomes analytically intractable, thus we employ the gradient-based scheme in Algorithm 1. The numerical experiments (see Fig. S1) show that the method converges consistently across all orientation patterns and produces solutions that agree with exhaustive or heuristic baselines, which indicates that the objective landscape remains smooth enough for reliable gradient-driven search.
The structural behaviors observed in the two-magnet case persist in the three-magnet configuration. More importantly, the aggregated results reveal clear structural tendencies that persist across all configurations. The first two magnets exhibit highly concentrated optimal placements near $L_{k_{0}}^{*}/L\approx 0.18$ and $L_{k_{1}}^{*}/L\approx 0.74$, reflecting a general preference for distributed magnetic sources that enrich the set of attainable torque directions. Orientation also plays a decisive role. Opposing pairs drive the optimizer to increase their spacing and shift them toward the proximal region, while like-oriented pairs naturally cluster toward one end to exploit larger moment arms. These consistent behaviors show that the interaction between magnet placement and orientation organizes the shape of the actuation space and dominates the resulting optimal designs.
V Optimization under General Fields
In this section, we extend the proposed structural optimization framework to dipole-generated fields and investigate the resulting equilibrium behavior, kinematic performance, and optimal embedded-magnet configurations.
V-A Numerical Optimization under Single Dipole Fields
TABLE I: Geometric and material properties
Parameters (Description) Value $\mbox{Unit}]$
Silicone tube $E$ (Young’s modulus) 2020 $\mathrm{MPa}]$
ODOD (Outside diameter) $1.2$ $\mathrm{mm}]$
IDID (Inside diameter) $0.8$ $\mathrm{mm}]$
$L$ (Length) $40$ $\mathrm{mm}]$
[…]
Micromagnets $B_{r}$ (Remanence) $1.2$ $\mathrm{T}]$
$L_{m}$ (Length) $5/3/2$ $\mathrm{mm}]$
$D_{m}$ (Diameter) 11 $\mathrm{mm}]$
$L_{g}$ (Minimal gap) $5/3/2$ $\mathrm{mm}]$
[…]
Actuating Magnet $M_{\mathrm{b}}$ (Dipole moment) $342.86$ $\mathrm{A\cdot m^{2}}]$
Figure 3: Experimental setup: A permanent magnet attached to a 6-DoF robot arm generates a spatially varying dipole field. The MeSCR is placed inside a transparent container, while a downward-facing camera records its deformation for subsequent shape extraction and Jacobian estimation. Insets show details of the MeSCR geometry (outer diameter $1.2\text{\,}\mathrm{m}\mathrm{m}$, embedded magnet length $5\text{\,}\mathrm{m}\mathrm{m}$) and an example of the extracted centerline.
Figure 4: Performance evaluation under three dipole-field task paths. a.) Illustration of the three task paths. b.)–f.) Normalized kinematic performances with respect to the movable magnet position $L_{k_{0}}/L$ with same and opposite directions of orientation: b.) $\mathcal{Z}_{\mathrm{vol}}$, c.) $\mathcal{Z}_{\mathrm{dis}}$, d.) $\mathcal{Z}_{\kappa}$, e.) $\mathcal{Z}_{\mathrm{vol}}/\mathcal{Z}_{\mathrm{dis}}$, and f.) $\mathcal{Z}_{\mathrm{vol}}/\mathcal{Z}_{\kappa}$. Solid curves correspond to numerical results for the three paths, and dashed lines indicate the optimal placements. Experimental data are plotted as discrete markers.
To study the effect of embedded magnet configurations under dipole fields, we evaluate the two principal orientation modes across representative task paths. The experimental setup is presented in Fig. 3, and the results comparison to simulations are summarized in Fig. 4, Fig. S2, and Table S1, which exhibit consistent trends that align with our analytical insights. For the same direction magnetization, all kinematic indices preserve their characteristic ordering from the uniform field analysis. The increasing actuation strength toward the distal end leads manipulability to prefer distal placement, while indices sensitive to curvature variation attain their maxima at intermediate spacing. These behaviors remain stable across task paths and magnet lengths, indicating that constructive torque accumulation dominates the optimization landscape.
Opposite direction magnetization produces a markedly different pattern. The interaction of opposing torques generates a cancellation zone that governs the deformation geometry, causing all performance measures to collapse toward a narrow proximal region. The near invariance across paths and magnet lengths confirms that this behavior is determined primarily by the intrinsic structure of the torque field rather than by task-specific sampling. The competing tendencies predicted in the uniform field case also persist: increasing magnet separation strengthens net actuation, whereas proximity favors smoother deformation. Under dipole actuation, however, the dominant role of cancellation compresses all optimal solutions toward a common proximal configuration. These results provide direct experimental validation that magnet orientation is the primary structural factor shaping the Jacobian spectrum and the geometry of the equilibrium manifold. Constructive or destructive torque interactions organize the qualitative behavior of all performance measures, while task paths and scaling parameters exert only secondary influence.
Figure 5: Optimal movable magnet position under varying values of the nondimensional parameter $\nu$. a.) Same direction magnetization. b.) Opposite direction magnetization.
We further examine the dependence of optimal placement on the nondimensional parameter
$\nu=\frac{4\pi EI}{N_{m}B_{r}M_{\mathrm{b}}d^{3}AL^{2}}.$
The results in Fig. 5 show that for the same direction magnetization, the optimal positions vary smoothly with $\nu$ and retain the established ordering. For opposite orientation, the optimal placements exhibit larger variation at small $\nu$ but converge rapidly toward boundary configurations as elasticity increases. The presence or absence of internal torque cancellation thus provides the underlying organizing principle for all indices across field strength, task path, and geometry.
V-B Numerical Optimization under General Fields
To evaluate whether the dipole–field behaviors observed in the two magnet case generalize to higher dimensional designs, we extend the analysis to MeSCRs with three embedded magnets, allowing two magnets to vary in both position and orientation. The experiments cover multiple dipole configurations, task paths, and magnetic strength parameters $(\nu_{1}\!-\!\nu_{5})$, with results summarized in Fig. S3. A clear and unified set of structural tendencies emerges across all settings. The placement hierarchy mirrors that of the two magnet case: constructive torque accumulation drives distal optimality for $\mathcal{Z}_{\mathrm{vol}}$, while indices sensitive to curvature modulation, such as $\mathcal{Z}_{\mathrm{dis}}$ and $\mathcal{Z}_{\kappa}$, consistently favor proximal configurations. Ratio based measures converge to stable mid–distal solutions, indicating path invariant behavior even under complex field variations. Orientation exhibits the same role identified previously. Aligned moments promote torque amplification and favor more concentrated arrangements, whereas partial spreading becomes advantageous when deformation diversity is required. The most consistent layouts arise from the combined indices, which naturally balance these competing trends. Together with the two magnet dipole–field results, these findings demonstrate that orientation dependent torque interactions and spacing driven actuation diversity form a unified geometric mechanism that persists across increased design dimensionality and external field complexity.
VI Conclusion
This work introduced a structural optimization framework that unifies equilibrium mechanics with kinematic performance for magnetically actuated soft continuum robots. By formulating magnetic loading as an actuation-dependent equilibrium manifold and evaluating Jacobian spectral indices on its pullback geometry, the approach enables physically consistent performance assessment across both analytical and numerically computed configurations. The results across uniform and dipole-generated fields reveal that the attainable motion of magnetic continuum structures is fundamentally organized by the interaction pattern of embedded magnets. Constructive torque interactions promote distal or mid–distal arrangements that strengthen net actuation, whereas opposing orientations create intrinsic cancellation regions that compress optimal designs toward proximal configurations. These tendencies persist across different actuation paths, field strengths, and magnet counts, indicating that orientation-driven torque geometry is a primary determinant of global kinematic behavior.
Certain limitations remain. The analytical development relies on small-deflection assumptions and a discrete PRB representation, which may reduce accuracy under strong fields or highly compliant materials. Magnet–magnet interactions, nonlinear constitutive effects, and dynamic behaviors are not explicitly modeled. Extending the approach to richer material descriptions, heterogeneous magnetization, and closed-loop actuation could broaden its practical utility. Further characterization of equilibrium-manifold geometry may deepen understanding of reachable sets, while integrating sensing, real-time control, or data-driven design tools could facilitate systematic structure actuation co-design for future magnetic soft continuum robots.